When heated, calcium carbonate decomposed part of the substance. At the same time 4.48 liters of gas were released. The mass of the solid residue was 33.2 g. This residue was added to 300 g of hydrochloric acid, taken in excess. Determine the mass fraction of salt in the solution.
1
Expert's answer
2020-06-29T07:46:30-0400

WHEN Heated Calcium Carbonate Decomposed gives CaO and CO2 .


THIS CaO Is Solid Residue And React with 2mol of HCl Gives CaCl2 (s) and H2O(l) .


CaCl2 is as a salt .


moles of CaO = 33.2 /56 . mol ..

left moles of CaO = 0 .



Moles of HCl reacted = 66.4 / 56 .mol ..

left moles of HCl = (300 / 65) - ( 66.4 /56 ) .



Moles of CaCl2 formed = 33.2 /56 = 0.592 mol..



Moles of H2O formed = 33.2 /56 mole ..



mass of salt formed =( 0.592 * 111 ) gm = 65.712 gm .

mass of H2O formed =( 33.2 *18 ) /56 . = 10.671 gm.

mass of HCl formed =( 7.034 * 36.5 ) gm .


Total Mass of solution left after reaction =( 65.712 + 10.67 + 256 .74 ) gm ...... = 333.124 gm ....


Mass fraction of salt in solution = mass of salt / mass of solution =

65.712 / 333.12 = 0.1972 . Answer .